OpenStudy (anonymous):

Simplify the complex fraction as much as possible. Leave your answer as a proper fraction or a mixed number. 1/2/ 1/3. Enter your answer as a reduced mixed number or as a whole number.

OpenStudy (anonymous):



OpenStudy (anonymous):

1/2 divided by 1/3 = 1/2 x 3/1 = 3/2 = 1 1/2